OpenStudy (goformit100):

@Nurali @whpalmer4 break the following function into intervals. y = | x -1 | + | x - 2 |

OpenStudy (anonymous):

That will give you two intervals. Same with \(|x-2|\). You just need to find where there is overlap.

OpenStudy (anonymous):

Well, \(x-1\) is either negative or not: \[ x-1<0 \quad \quad 0\leq x-1 \]

OpenStudy (goformit100):

What if the given function is given to be differentiate, what to do explain it briefly ....

OpenStudy (anonymous):

you need to differentiate the function? use chain rule.

OpenStudy (anonymous):

\[ |x| = \sqrt{x^2} \]

OpenStudy (goformit100):

How please Explain it stepwise and deeply

OpenStudy (anonymous):

Okay, solve \(x-1<0\). Can you do it?

OpenStudy (anonymous):

then solve \(x-2<0\).

OpenStudy (anonymous):

Then you have \(x<a\), the intervals \(a\leq x<b\), and \(b\leq x\)

OpenStudy (goformit100):

not getting it

OpenStudy (anonymous):

\[ |x-a| = \begin{cases}-(x-a)&x<a\\ x-a&a\leq x\end{cases} \]

OpenStudy (anonymous):

\[ (f\circ g)(x) = |x|= \sqrt{x^2}\\ f(x) = \sqrt{x} \quad f'(x)=\frac{1}{2\sqrt x}\\ g(x) = x^2\quad g'(x) = 2x \\ (f\circ g)'(x) = (f'\circ g)(x) \cdot g'(x) = \frac{1}{2\sqrt{x^2}}(2x) = \frac{x}{\sqrt{x^2}}=\frac{x}{|x|} \]

OpenStudy (anonymous):

In short \[ |x|' = \frac{x}{|x|} \]

OpenStudy (anonymous):

Is that enough for you?

OpenStudy (goformit100):

Yes Sir

OpenStudy (kainui):

@goformit100 I suggest you sit down and try to graph specific points on this graph yourself. See what it looks like. Sure it will take a little time, but you'll understand through self discovery. Remember, the turning points of an absolute value are when it equals 0 to go between negative and positive.

OpenStudy (goformit100):

Thanks for the suggestion Sir @Kainui